OpenStudy (anonymous):

rationalise the denominator of 4+√5 / √5

OpenStudy (goformit100):

Multiply √5 in numerator and denominator

OpenStudy (goformit100):

@Miracrown ma'am can teach you the best.

OpenStudy (anonymous):

i have done that but I keep getting the answer wrong , here is my workings out for the question 4 + √5 * √5 = 4+ √25 = 4+5 and then for the denominator it would be 5 , but the answer in the textbook says 4+ √5 + 5 / 5

OpenStudy (anonymous):

Multiply by \[\frac{ 4+\sqrt{5} }{ \sqrt{5} }\times \frac{ \sqrt{5} }{ \sqrt{5}}\]

OpenStudy (anonymous):

You have to distribute sq root of 5 to both 4 and saquare root of 5.\[4 + \sqrt{5} \times \sqrt{5} = 4\sqrt{5} + \sqrt{25} = 4\sqrt{5}+5\]

OpenStudy (skullpatrol):

Please use grouping symbols.

OpenStudy (skullpatrol):

There is a good reason the distributive property is written with grouping symbols as a(b+c) not a*b+c

OpenStudy (anonymous):

What does rationalize mean? What subject is this part of? (Algebra?) Sorry, I'm not sure of the answer but I'm interested in learning about this. :) Good luck with your homework~